### v1

@@ Abstract
Peer prediction mechanisms incentivize truthful reporting of private signals in settings where ground truth is unavailable. A truthful mechanism, which ensures that truth-telling maximizes the expected scores, can elicit honest information while assuming agents’ utilities are linear functions of their scores. However, in many practical scenarios, non-linear payment rules are preferred, or agents’ utilities are inherently non-linear. For example, students in peer grading may care about letter grades rather than the numerical scores computed by the mechanism.
In this paper, we introduce the concept of stochastically dominant truthfulness (SD-truthfulness), which requires that, in equilibrium, the score distribution of any non-truthful strategy is first-order stochastically dominated by that of truth-telling. An SD-truthful mechanism can elicit honest information as long as agents’ utilities are monotone increasing functions of their scores. To the best of our knowledge, no existing peer prediction mechanism naturally satisfies this criterion without strong assumptions.
We first identify a common rounding reduction that can transform any truthful mechanism into an SD-truthful mechanism by normalizing scores into probabilities and awarding binary scores via a lottery. However, this reduction often destroys sensitivity, an important property concerning the fairness and efficiency of a mechanism. We overcome this limitation by showing how a more nuanced application of rounding can partially recover the sensitivity of the original mechanism. Furthermore, we propose a novel enforced agreement (EA) mechanism that achieves SD-truthfulness in binary-signal settings under mild assumptions. Our empirical results show that EA achieves the highest sensitivity among all known SD-truthful mechanisms. Unfortunately, this approach does not generalize beyond the binary-signal setting, highlighting a key challenge in obtaining highly sensitive SD-truthfulness in general settings.
@@ 1 Introduction
Information elicitation studies how to design mechanisms that incentivize honest and high-effort human feedback. In this framework, a principal seeks reliable responses regarding a set of tasks from strategic agents who are motivated by the designed rewards. For example, AI companies collect human preferences to align language models via RLHF (Ouyang et al., 2022), and MOOCs use peer grading to evaluate complex assignments at scale (Piech et al., 2013). A key challenge is designing scoring mechanisms that accurately assess the informativeness of agents’ reports.
Peer prediction provides an elegant solution for scenarios where the quality of collected reports cannot be directly verified against ground truth (Miller et al., 2005). This method is particularly valuable when obtaining ground truth is costly or infeasible, such as in subjective tasks. The core idea is to evaluate each agent based on the correlation between their reports and those of a randomly chosen peer. The scoring mechanism is designed to be truthful so that everyone reporting their information truthfully forms a Bayesian Nash equilibrium, ensuring that truth-telling maximizes each agent’s expected score.
A key assumption underlying the effectiveness of existing peer prediction mechanisms is that agents’ utility is a linear function of their score. This assumption is reasonable in some cases, such as when agents are solely motivated by monetary payments, which can be straightforwardly designed as a linear transformation of the peer prediction scores. However, practical implementations often favor non-linear transformations, such as tournaments, to map scores to payments due to their improved budgetary efficiency and flexibility (Zhang and Schoenebeck, 2023a). Moreover, agents’ intrinsic utility functions are sometimes inherently non-linear. For instance, student graders care more about their final letter grades (e.g., A, B, C) than the real-valued numerical scores assigned by the mechanism. When agents’ utilities are non-linear functions of their scores, truth-telling is no longer guaranteed to be an equilibrium strategy.
Motivated by this key limitation, this paper proposes the concept of stochastic dominance-truthful (SD-truthful) peer prediction mechanisms. Under this stronger truthful guarantee, we aim to design mechanisms in which an agent’s score distribution, when everyone reports truthfully, first-order stochastically dominates the score distribution if they unilaterally deviate to any other strategy. As a result, an SD-truthful mechanism incentivizes truth-telling for agents with any utility function that is monotone-increasing with the score.
@@ 1.1 Our Results
Our first observation is that no existing peer prediction mechanism is naturally SD-truthful under general information structures. We investigate three widely studied mechanisms that compute scores as the average of multiple simple discrete scores, including Output Agreement (OA), Peer Truth Serum (PTS) (Faltings et al., 2017), and Correlated Agreement (CA) (Shnayder et al., 2016). Mechanisms with complex scoring rules — such as those that rely on continuous-valued unbiased estimates of mutual information (Kong and Schoenebeck, 2019; Kong, 2020; Schoenebeck and Yu, 2020) — are unlikely to satisfy SD-truthfulness.
However, even these three simple mechanisms fail to achieve SD-truthfulness without imposing strong assumptions on the underlying information structure. In particular, OA is truthful and SD-truthful only when signals are self-dominating, meaning that if an agent observes a signal $\sigma$, their belief that another agent also observes $\sigma$ is always the highest. While PTS and CA remain truthful under more general information structures, they are not naturally SD-truthful.
@@
Our second observation is that achieving SD-truthfulness is easy in principle but challenging in practice. Given any truthful mechanism $\mathcal{M}$ with bounded scores $S^{\mathcal{M}}\in[S_{\inf},S_{\sup}]$, we can define a probability by normalizing the score: $\lambda^{\mathcal{M}}=\frac{S^{\mathcal{M}}-S_{\inf}}{S_{\sup}-S_{\inf}}\in[0,1]$. We then score agents via a binary lottery: $\tilde{S}^{\mathcal{M}}=1$ with probability $\lambda^{\mathcal{M}}$ and 0 otherwise. However, although this approach, called the direct rounding reduction, trivially satisfies SD-truthfulness, it greatly reduces sensitivity — an important property related to the efficiency and ex-post fairness of a mechanism. At a high level, sensitivity measures how much a mechanism’s score responds to changes in the information contained in an agent’s reports, normalized by the standard deviation of the scores.
@@
To overcome this limitation, we propose the partition rounding reduction, which helps partially recover the sensitivity of the original mechanism. The key idea is to divide the $n$ tasks into $K$ disjoint partitions and apply the original mechanism separately within each partition to compute an individual score $S^{\mathcal{M}}_{i}$. The final score for an agent is then the average of $K$ i.i.d. individual scores after applying the direct rounding if necessary. Intuitively, this approach improves sensitivity by reducing score variance through the aggregation of multiple independent individual scores.
@@
However, the partition rounding process cannot fully recover the sensitivity of the original mechanisms, implying additional room for more sensitive peer prediction mechanisms. We introduce a novel mechanism called enforcement agreement (EA) that is shown to be the most sensitive SD-truthful mechanism in the binary-signal setting. EA aims to replicate the key advantage of OA — its high sensitivity — while relaxing the strong assumptions on the information structure required for SD-truthfulness.
@@
EA works by enforcing a pre-determined empirical distribution of signals. In the binary setting, if Alice reports 0 on $m_{0}>n_{0}$ questions (where $n_{0}$ is the enforced count), the mechanism randomly flips $m_{0}-n_{0}$ of them to 1. This trick removes agents’ incentive to over-report majority signals — a strategy that undermines the truthful properties of OA. However, EA is only truthful but not SD-truthful when signals are non-binary, highlighting an important future challenge in designing high-sensitivity SD-truthful mechanisms for more complex settings.
@@
Lastly, our empirical results reinforce and complement our theoretical insights. In the binary-signal setting, we show that even a prior-free implementation of EA — simply enforcing a uniform distribution — outperforms other SD-truthful mechanisms in most settings, including those that require prior knowledge. For non-binary-signal settings, the partition rounding reduction of PTS emerges as the most sensitive SD-truthful mechanism. As a robustness check, we validate our findings by showing that EA consistently achieves the highest budgetary efficiency, a practical metric of peer prediction mechanisms that quantifies the minimum budget required to elicit a desired level of effort (Zhang and Schoenebeck, 2023a).
@@
Table 1 summarizes a comparison of SD-truthful (implementations of) peer prediction mechanisms with respect to the assumptions they require and their sensitivities. [^1] Matching Agreement (MA) is an extension of Correlated Agreement (CA), and therefore shares similar requirements and properties.
@@
Mechanisms Requirements for SD-truthfulness Output Agreement (OA) Self-dominating signal Enforced Agreement (EA) (This paper) Binary and self-predicting signal Peer Truth Serum (PTS) (Faltings et al., 2017) Self-predicting signal Correlated Agreement (CA) (Shnayder et al., 2016) Any signal Matching Agreement (MA) (Zhang and Schoenebeck, 2023b) Any signal
@@ 2 Related Work
Miller et al. (2005) first propose the idea of peer prediction — using a peer’s report to validate an agent’s report via proper scoring rules. Their mechanism is not detail-free, meaning that the mechanism requires prior knowledge of the information structure to obtain truthfulness. There are two ideas to mitigate this limitation: 1) the single-task mechanism, which elicits the information structure directly from agents (Prelec, 2004); and 2) the multi-task mechanism, which assumes each agent responds to multiple i.i.d. tasks but only requires eliciting signals rather than predictions (Dasgupta and Ghosh, 2013).
Our study is primarily relevant to the multi-task peer prediction literature. Connecting to our problem, we categorize multi-task peer prediction mechanisms into two types: peer agreement mechanisms (Dasgupta and Ghosh, 2013; Faltings et al., 2017; Shnayder et al., 2016; Zhang and Schoenebeck, 2023b; Agarwal et al., 2017) and mutual information mechanisms (Kong and Schoenebeck, 2019; Kong, 2020; Schoenebeck and Yu, 2020; Schoenebeck et al., 2021; Kong, 2024; Lu et al., 2024).
The score of a peer agreement mechanism can be viewed as an average of multiple individual scores, where each individual score usually takes values from a small finite space, e.g. $\{0,1\}$. For example, the output agreement (OA) mechanism scores an agent based on the average number of questions that they agree with another agent. The first multi-task peer prediction mechanism proposed by Dasgupta and Ghosh (2013) lies in this category, which can achieve truthfulness in a binary-signal setting. [^2] In fact, their mechanism is strongly truthful, which guarantees the truth-telling equilibrium uniquely maximizes each agent’s expected score among all strategy profiles. The correlated agreement (CA) mechanism generalizes their idea to any finite-signal setting (Shnayder et al., 2016). Later, Kong and Schoenebeck (2019) show that the payment of CA is actually an estimate of total variation distance mutual information. The peer truth serum (PTS) mechanism can be viewed as a generalization of OA, generalizing the self-dominating assumption to self-prediction, which basically requires signals to be positively correlated in the binary setting. In addition, there are two follow-ups of CA aiming to handle heterogeneous agents (Agarwal et al., 2017) and more general cheating strategies (Zhang and Schoenebeck, 2023b).
@@
In this paper, we primarily focus on peer agreement mechanisms because stochastic dominance-truthfulness implies that the score of the mechanism needs to take a simple form. Intuitively, this is because a complex scoring mechanism provides more opportunities for a cheating strategy to obtain one of the highest scores with a larger probability than truth-telling, which breaks SD-truthfulness.
Therefore, mutual information mechanisms that score each agent by estimating the mutual information between their reports and their peers’ reports are unlikely to obtain SD-truthfulness, as these estimations are usually complicated. In contrast, the determinant mutual information mechanism scores an agent based on the determinant of the estimated joint distribution between two agents’ reports (Kong, 2020); and the score of the $\Phi$-pairing mechanism needs to go through a complex machine learning process. We thus consider more detailed discussions of these mechanisms beyond the scope of this paper.
@@
We further note that the idea of rounding peer prediction scores into a binary lottery was previously proposed by Miller et al. (2005) to address risk-averse agents. Thus, at least the direct rounding reduction discussed in Section 5 is not novel. What is new in our work is the comparative analysis of various rounding reductions based on sensitivity, allowing us to identify a more effective rounding approach. Furthermore, we formally prove that this idea extends beyond risk-averse agents and can be applied to any agents with increasing utility functions.
@@ 3 Model
We consider the multi-task peer prediction setting where two agents, Alice and Bob, answer the same $n$ i.i.d. questions. For each question $j$, agents receive discrete signals $X_{a,j},X_{b,j}\in\Sigma=\{0,\ldots,c-1\}$, drawn from a fixed joint distribution $\Pr(X_{a},X_{b})$. With more than two agents, Bob is chosen randomly as Alice’s peer for scoring.
@@ 3.1 Truthful Guarantees
Existing peer prediction mechanisms usually focus on the expected scores, guaranteeing that deviations from truth-telling will only reduce the expected scores. In other words, truth-telling is a Bayesian Nash Equilibrium. Because we focus on truth-telling under the equilibrium, we assume Bob is always truthfulg. This means that the score of a peer prediction mechanism reduces to a function of Alice’s strategy $\theta$, denoted as $S^{\mathcal{M}}(\theta)$.
Definition 1.
A peer prediction mechanism is truthful if the expected score $\mathbb{E}[S(\tau)]\geq\mathbb{E}[S(\theta)]$ for any $\theta\in\Theta$.
A truthful mechanism guarantees that if agents’ utilities are linear functions of the scores, truth-telling is an equilibrium — any deviation from truth-telling will result in a smaller expected utility. Previous works have explored stronger notions of truthfulness, such as: informed truthfulness (Shnayder et al., 2016), which guarantees that truth-telling provides each agent with the highest expected score and yields a strictly higher score whenever an agent deviates to an uninformative strategy; and strong truthfulness (Dasgupta and Ghosh, 2013), which ensures that truth-telling uniquely maximizes every agent’s expected score. However, these concepts deal with the equilibrium selection among multiple equilibria, whereas in this paper, we are concerned with the properties of the truth-telling equilibrium.
However, as we argued earlier, the linear-utility assumption is often undesirable and, in some cases, unrealistic. In this paper, we want to design mechanisms that are truthful for a much broader and more reasonable class of utility functions — specifically, any utility function that is increasing in score. This motivation naturally leads us to the concept of stochastic dominance.
Definition 2.
A peer prediciton mechanism is stochastic dominance-truthful (SD-truthful) if the distribution of $S(\tau)$ first-order stochastic dominates (FOSD) the distribution of $S(\theta)$ for any $\theta\in\Theta$, i.e. $\Pr(S(\tau)\geq t)\geq\Pr(S(\theta)\geq t)$ for any $t\in\mathbb{R}$.
A standard result in microeconomics suggests that $S(\tau)$ FOSD $S(\theta)$ if and only if $\mathbb{E}[u(S(\tau))]\geq\mathbb{E}[u(S(\theta))]$ holds for every increasing utility function $u$ (Hadar and Russell, 1969). Since linear utility functions are a special case of increasing utility functions, it follows that every SD-truthful mechanism is also truthful.

### v2

@@ Abstract
Eliciting reliable human feedback is essential for many machine learning tasks, such as learning from noisy labels and aligning AI systems with human preferences. Peer prediction mechanisms incentivize truthful reporting without ground truth verification by scoring agents based on correlations with peers. Traditional mechanisms, which ensure that truth-telling maximizes the expected scores in equilibrium, can elicit honest information while assuming agents’ utilities are linear functions of their scores. However, in practice, non-linear payment rules are usually preferred, or agents’ utilities are inherently non-linear.
We propose stochastically dominant truthfulness (SD-truthfulness) as a stronger guarantee: the score distribution of truth-telling stochastically dominates all other strategies, incentivizing truthful reporting for a wide range of monotone utility functions. Our first observation is that no existing peer prediction mechanism naturally satisfies this criterion without strong assumptions. A simple solution—rounding scores into binary lotteries—can enforce SD-truthfulness, but often degrades sensitivity, a key property related to fairness and statistical efficiency. We demonstrate how a more careful application of rounding can better preserve sensitivity. Furthermore, we introduce a new enforced agreement (EA) mechanism that is theoretically guaranteed to be SD-truthful in binary-signal settings and, under mild assumptions, empirically achieves the highest sensitivity among all known SD-truthful mechanisms.
@@ 1 Introduction
Information elicitation studies how to design mechanisms that incentivize honest and high-effort human feedback. In this framework, a principal seeks reliable responses regarding a set of tasks from strategic agents who are motivated by the designed rewards. For example, AI companies collect human preferences to align language models via RLHF (Ouyang et al., 2022), and MOOCs use peer grading to evaluate complex assignments at scale Piech et al. (2013). A key challenge is designing scoring mechanisms that accurately assess the informativeness of agents’ reports.
Peer prediction provides an elegant solution for settings where ground truth is unavailable or costly to obtain, such as in subjective tasks (Miller et al., 2005). Instead of relying on external validation, peer prediction scores agents based on the correlation between their reports and those of their peers. The scoring mechanism is designed to be truthful so that everyone reporting truthfully forms a Bayesian Nash equilibrium, ensuring that truth-telling maximizes each agent’s expected score.
A key assumption underlying the effectiveness of existing peer prediction mechanisms is that agents’ utility is a linear function of their score. This assumption is reasonable in some cases, such as when agents are solely motivated by monetary payments, which can be straightforwardly designed as a linear transformation of the peer prediction scores. However, in practice, non-linear payment rules such as threshold bonuses that reward workers only when performance exceeds a cutoff are often preferred for their budget efficiency and flexibility (Zhang and Schoenebeck, 2023a). Moreover, agents’ intrinsic utility functions are sometimes inherently non-linear. For instance, student graders care more about their final letter grades (e.g., A, B, C) than the real-valued numerical scores assigned by the mechanism. When agents’ utilities are non-linear functions of their scores, truth-telling is no longer guaranteed to be an equilibrium strategy.
Motivated by this limitation, we introduce the concept of stochastic dominance-truthful (SD-truthful) peer prediction mechanisms. This stronger notion of truthfulness requires that an agent’s score distribution, when everyone reports truthfully, first-order stochastically dominates the score distribution under any unilateral deviation. As a result, SD-truthful mechanisms incentivize truth-telling for agents with any utility function that is monotone increasing with the score.
@@ 1.1 Our Results
Our first observation is that no existing peer prediction mechanism is naturally SD-truthful under general information structures, i.e., under arbitrary correlations between agents’ signals. We examine three widely studied mechanisms that compute scores as the average of multiple simple discrete scores, including Output Agreement (OA), Peer Truth Serum (PTS) (Faltings et al., 2017), and Correlated Agreement (CA) (Shnayder et al., 2016). Mechanisms with complex scoring rules—such as those based on continuous-valued estimates of mutual information Kong and Schoenebeck (2019); Kong (2020); Schoenebeck and Yu (2020)—are unlikely to satisfy SD-truthfulness. Yet even these simple mechanisms fail to guarantee SD-truthfulness without strong assumptions on the underlying information structure.
@@
Our second observation is that achieving SD-truthfulness is easy in principle but challenging to do well in practice. Given any truthful mechanism $\mathcal{M}$ with bounded scores $S^{\mathcal{M}}\in[S_{\inf},S_{\sup}]$, we can define a probability by normalizing the score: $\lambda^{\mathcal{M}}=\frac{S^{\mathcal{M}}-S_{\inf}}{S_{\sup}-S_{\inf}}\in[0,1]$. We then score agents via a binary lottery: $\tilde{S}^{\mathcal{M}}=1$ with probability $\lambda^{\mathcal{M}}$ and 0 otherwise. Although this approach, called the direct rounding reduction, trivially satisfies SD-truthfulness, it greatly reduces sensitivity—an important property related to the efficiency and ex-post fairness of a mechanism. At a high level, sensitivity measures how much a mechanism’s score responds to changes in the information contained in an agent’s reports, normalized by the standard deviation of the scores.
@@
To overcome this limitation, we propose the partition rounding reduction, which helps partially recover the sensitivity of the original mechanism. The idea is to divide the questions into $K$ disjoint subsets and apply the original mechanism separately within each subset to compute an individual score. The final score for an agent is then the average of $K$ i.i.d. individual scores after applying direct rounding. Intuitively, this approach improves sensitivity by reducing variance through the aggregation of multiple independent scores.
@@
However, the partition rounding process cannot fully recover the sensitivity of the original mechanisms, implying additional room for more sensitive peer prediction mechanisms. We introduce a novel mechanism called enforced agreement (EA) that is shown to be the most sensitive SD-truthful mechanism in the binary-signal setting. EA aims to replicate the key advantage of OA—its high sensitivity—while relaxing the strong assumptions on the information structure required for SD-truthfulness.
@@
EA works by enforcing a pre-determined empirical distribution of signals. In the binary setting, if Alice reports 0 on $m_{0}>n_{0}$ questions (where $n_{0}$ is the enforced count), the mechanism randomly flips $m_{0}-n_{0}$ of them to 1. This trick removes agents’ incentive to over-report majority signals, a strategy that undermines the truthful properties of OA. However, EA is only truthful but not SD-truthful when signals are non-binary, highlighting an important future challenge in designing high-sensitivity SD-truthful mechanisms for more complex settings.
@@
Lastly, our empirical results reinforce and complement our theoretical insights. In the binary-signal setting, we show that even a prior-free implementation of EA—simply enforcing a uniform distribution—outperforms other SD-truthful mechanisms in most settings, including those that require prior knowledge. For non-binary-signal settings, the partition rounding reduction of PTS emerges as the most sensitive SD-truthful mechanism.
@@
Table 1 summarizes a comparison of SD-truthful (implementations of) peer prediction mechanisms with respect to the assumptions they require and their sensitivities. [^1] MA is an extension of CA, and therefore shares similar requirements and properties.
@@
Mechanisms Requirements for SD-truthfulness Output Agreement (OA) Self-dominating signal Enforced Agreement (EA) [This paper] Binary and self-predicting signal Peer Truth Serum (PTS) (Faltings et al., 2017) Self-predicting signal Correlated Agreement (CA)(Shnayder et al., 2016) Any signal Matching Agreement (MA) (Zhang and Schoenebeck, 2023b) Any signal
@@ 2 Related Work
The idea of peer prediction was proposed by Miller et al. (2005). As the first discussion, their mechanism is not detail-free, i.e. the mechanism requires prior knowledge of the information structure (how agents’ signals are correlated) to obtain truthfulness. There are two ideas to mitigate this limitation: (i) the single-task mechanism, which elicits the information structure directly from agents (Prelec, 2004); and (ii) the multi-task mechanism, which assumes each agent responds to multiple i.i.d. tasks but only requires eliciting signals rather than predictions (Dasgupta and Ghosh, 2013).
@@
Our study is primarily relevant to the multi-task setting. We categorize multi-task peer prediction mechanisms into two types: peer agreement mechanisms Dasgupta and Ghosh (2013); Faltings et al. (2017); Shnayder et al. (2016); Zhang and Schoenebeck (2023b); Agarwal et al. (2017) and mutual information mechanisms Kong and Schoenebeck (2019); Kong (2020); Schoenebeck and Yu (2020); Schoenebeck et al. (2021); Kong (2024). In this paper, we primarily focus on peer agreement mechanisms, as SD-truthfulness requires the mechanism’s score to take a relatively simple form. Intuitively, this is because a complex scoring mechanism provides more opportunities for a cheating strategy to obtain one of the highest scores with a larger probability than truth-telling, which breaks SD-truthfulness. In Appendix A, we provide a more detailed discussion on why several classic mutual information mechanisms are not SD-truthful.
The score of a peer agreement mechanism can be viewed as an average of multiple individual scores, where each individual score takes values from a small finite space, e.g. $\{0,1\}$. For example, output agreement (OA) scores an agent based on the average number of questions that they agree with another agent. The first multi-task peer prediction mechanism lies in this category, which can achieve truthfulness in a binary-signal setting (Dasgupta and Ghosh, 2013). The correlated agreement (CA) mechanism generalizes their idea to any finite-signal setting (Shnayder et al., 2016). The peer truth serum (PTS) mechanism is a generalization of OA, which relaxes the self-dominating assumption to self-prediction. In addition, two follow-ups of CA aim to handle heterogeneous agents (Agarwal et al., 2017) and more general cheating strategies (Zhang and Schoenebeck, 2023b).
@@
Moreover, we note that the idea of rounding peer prediction scores into a binary lottery (see Section 5) was previously proposed by Miller et al. (2005) to address risk-averse agents. What is new in our work is the comparative analysis of various rounding reductions based on sensitivity, allowing us to identify a more effective rounding approach. Furthermore, we formally prove that this idea extends beyond risk-averse agents and can be applied to any agents with increasing utility functions.
@@ 3 Model
We consider the multi-task peer prediction setting where two agents, Alice and Bob, answer the same $n$ i.i.d. questions. For each question $j$, the agents receive discrete signals $X_{a,j},X_{b,j}\in\Sigma=\{0,\ldots,c-1\}$, drawn from a fixed joint distribution $\Pr(X_{a},X_{b})$. In cases with more than two agents, we can reduce to the case of two agents by, for any particular agent “Alice”, choosing a random peer agent “Bob” for scoring.
@@ 3.1 Truthful Guarantees
Existing peer prediction mechanisms focus on expected scores, guaranteeing that any deviation from truth-telling will only reduce the expected score. In this sense, truth-telling forms a Bayesian Nash Equilibrium. Since our analysis centers on equilibrium behavior, we assume Bob always tells the truth. Under this assumption, the score assigned by the mechanism reduces to a function of Alice’s strategy $\theta$, written as $S^{\mathcal{M}}(\theta)$.
Definition 3.1.
A peer prediction mechanism is truthful if $\mathbb{E}[S(\tau)]\geq\mathbb{E}[S(\theta)]$ for any $\theta\in\Theta$.
A truthful mechanism guarantees that if agents have linear utility in their scores, truth-telling is a Bayesian Nash Equilibrium—any deviation yields lower expected utility. However, as we argued earlier, assuming linear utility is often undesirable and, in some cases, unrealistic. Instead, we aim to design truthful mechanisms for a broader and more reasonable class of utility functions—those that are simply increasing in score. This motivation naturally leads to the concept of stochastic dominance.
Definition 3.2.
A peer prediction mechanism is stochastic dominance-truthful (SD-truthful) if the distribution of $S(\tau)$ first-order stochastic dominates (FOSD) the distribution of $S(\theta)$ for any $\theta\in\Theta$, i.e. $\Pr(S(\tau)\geq t)\geq\Pr(S(\theta)\geq t)$ for any $t\in\mathbb{R}$.
A standard result in microeconomic theory suggests that $S(\tau)$ FOSD $S(\theta)$ if and only if $\mathbb{E}[u(S(\tau))]\geq\mathbb{E}[u(S(\theta))]$ for every increasing utility function $u$ (Hadar and Russell, 1969). Since linear utility functions are also increasing, every SD-truthful mechanism is also truthful.